er (15-1)(14) 2 howto
Answer #1

Sample size = n = 15

Sample mean = \bar{x} = 25

Standard deviation = s = 32

We have to construct 98% confidence interval.

Formula is

(\bar{x}-E , \bar{x}+E)

Here E is a margin of error.

E=

Degrees of freedom = n - 1 = 15 - 1 = 14

Level of significance = 0.02

tc = 2.624   ( Using t table)

E = \frac{2.624*3.2}{\sqrt{15}}=2.1680

So confidence interval is ( 25 - 2.1680 , 25 + 2.1680) = > ( 22.8320 , 27.1680)